Two electric charges one 20 times as strong as the other, exert a force 2.45×10-2 on each other. When they are placed 10 cm apart in air, find the magnitude of each charge.
1
Expert's answer
2020-06-03T12:10:32-0400

Explanations & Calculations

  • This is quiet straightforward as it's all about applying the Coulomb's equation for the situation.
  • "F = \frac{1}{4\pi\epsilon_0}\Big(\frac{q_1q_2}{r^2}\Big) =k\Big(\frac{q_1q_2}{r^2}\Big)"
  • k = 9 "\small \times" 109 Nm2C-2 in air.


  • If the charge of one charge is q then,

"\qquad\qquad
\begin{aligned}
\small 2.45 \times 10^{-2}N &= \small 9\times10^9\Bigg(\frac{q\times20q}{0.1^2m^2}\Bigg)\\
\small q^2 &= \small 1.3611\times 10^{-15}C^2\\
\small |q| &= \small 3.689\times10^{-8}C \\
&= \small \bold{0.0369\,\mu C}\\
\\
\small |20q| &= \small 20\times 36.89 = \small \bold{0.738 \,\mu C }
\end{aligned}"